### v2

@@ B.3 Generalization and Integration with Existing Methods
Table 15: Comparison of Ranksim, Ordinal Entropy, and AdaPRL-integrated methods on the KuaiRec, AgeDB, and Jigsaw datasets. Rel.Imp denoted the relative improvement over the corresponding metrics.
In this section, we evaluated the generalizability of AdaPRL and its compatibility with previously proposed methods. As shown in Table 1, AdaPRL demonstrated a significant improvement over the traditional L2 loss function across the majority of regression datasets. However, when comparing AdaPRL employed directly as $L_{reg}$ with the Ranksim and Ordinal Entropy methods, AdaPRL performed sightly inferiorly on certain datasets such as KuaiRec, AgeDB and Jigsaw. Consequently, it was essential to explore whether integrating Ranksim and Ordinal Entropy with AdaPRL could further boost the performance. Therefore, we designed new experiments on these three datasets by combining AdaPRL with both Ranksim and Ordinal Entropy, aiming to assess the resultant performance.
Specifically, the hyperparameters of both Ranksim and Ordinal Entropy method were optimized in each dataset via a coarse-grained grid search, and the weights were searched across {0.05,0.1,0.2,0.4,0.8,1.6}. Secondly, as we described in algorithm1, we computed and added an alpha-weighted $L_{CPRL}$ to the loss function of Ranksim and Ordinal Entropy with corresponding optimized hyperparameter. Then a coarse-grained grid search was performed on $\alpha$ of AdaPRL. The grid search and other settings were kept the same as the early experiment details.
The results, presented in Table 15, indicated that the integration of AdaPRL with Ranksim and Ordinal Entropy led to additional performance improvements across all three datasets. This underscored the versatility of AdaPRL in regression tasks and demonstrated its capability to be seamlessly incorporated into various regression frameworks.
B.4 Robustness to label noise in AdaPRL
Label noise is a common problem in real-world datasets, where mislabeling can occur due to data ambiguity, human error, automated labeling process, or intrinsic randomness of data, which can significantly degrade the performance of deep neural networks because of overfitting. As we had mentioned before, AdaPRL could theoretically alleviate aleatoric noise of the data because of the uncertainty estimation module. To prove the theory, we evaluated our method against standard baseline models using the KuaiRec dataset under varying levels of Gaussian noise for labels. The noise was introduced into the labels at different levels, where the label noise level $k$ signified that Gaussian noise with zero mean and a standard deviation of $0.2k\times\mathrm{std}(y)$ was added to the labels. $\mathrm{std}(y)$ denoted the standard deviation of the original labels here. This setup allowed us to assess the robustness of our method in the presence of different degrees of label noise. The experiment was repeated five times and the average of mean squared error (MSE) of test data was reported. The result was presented in Figure 5, AdaPRL consistently outperformed the baseline, exhibiting significantly less performance degradation as the proportion of noisy labels increased. Furthermore, as the level of label noise increased, our proposed method exhibited greater relative improvements in the MSE metric compared to the traditional MSE loss. This indicated that our method effectively mitigated the impact of label noise, resulting in more reliable model performance even when training data labels were imperfect.
@@ B.4 Robustness to label noise in AdaPRL
B.5 Robustness to unforeseen data corruption in AdaPRL
Ensuring that deep neural networks maintain performance in the presence of unforeseen data corruptions in the inferring stage is essential for their deployment in real-world applications. Models often encounter various unexpected corruptions or inconsistencies, such as missing data, noise from data collection, and inconsistencies of online and offline features, which could significantly degrade online performance. To simulate the problems, we chose the Movielens-1M and KuaiRec datasets as representative, and the evaluation protocol was as follows. Firstly, we randomly selected 20% of the columns to corrupt. Secondly, for each corruption level $k$, selected $10k\%$ of the validation and test samples. Then, randomly shuffled the values within the selected samples for each column. Finally, trained the model with the original training data and evaluated it on the corrupted test set. Note that in this experiment our model was trained solely on uncorrupted data, but was evaluated on corrupted test sets. Early stoppings were applied on the original validation data and a coarse-grained grid search over alpha was conducted as mentioned above. The experiment was repeated five times and the average of mean squared error (MSE) of test data was reported.
@@ B.5 Robustness to unforeseen data corruption in AdaPRL
As was shown in Figure 6, our method consistently outperformed the baseline, exhibiting better performance compared to the original L2 loss. Moreover, our method consistently achieved greater relative improvements in mean squared error (MSE) metrics compared to the L2 loss, even as the severity of data corruption increased, which we believed demonstrated the model’s improved resilience to unforeseen data corruptions.
B.6 Resilience to reduced training data in AdaPRL
Although the success of deep learning models is heavily dependent on the scale of training datasets, acquiring large quantities of labeled data is often impractical due to resource constraints. In many real-world situations, models must be effective even with limited training data. In this section, we conducted experiments by creating random subsets of the KuaiRec dataset with varying fractions to simulate scenarios with limited data. Our method was compared against a standard L2 loss baseline under the same conditions. The experiment was repeated five times and the average of mean squared error (MSE) of test data was reported.
@@ B.6 Resilience to reduced training data in AdaPRL
As was shown in Figure 7, our model demonstrated superior performance with significantly less degradation as the number of training samples decreased. This indicated that AdaPRL was more effective in settings where training data was limited, making it a viable solution for applications with constrained data availability. Another conclusion could be drawn that AdaPRL outperformed L2 loss in all cases, and exhibited the highest relative improvement when the number of training data was at a moderate level.
@@ B.7 Visualization of Time Series Forecasting Prediction Results
To facilitate an instinct comparison among different backbone forecasting models with or without AdaPRL, we presented prediction visualizations for two representative datasets in Figure 8. The figure illustrated the prediction results produced by the following models: iTransformerLiu et al. (2024) and MinusformerLiang et al. (2024) with a prediction length of 336. Upon visual inspection of the prediction results across these datasets, both iTransformer and Minusformer trained with AdaPRL demonstrated better accuracy in capturing future series variations, exhibiting superior performance in long-term forecasting compared to using MSE loss, which was consistent with the previous experimental result.
@@ B.8 Visualization of Time Series Forecasting Uncertainty Estimation
Moreover, by integrating a deep probabilistic network with an arbitrary backbone network, our proposed AdaPRL is capable of generating not only point estimates for each timestamp in long-term time series forecasting but also predicting the associated uncertainties. Figure 9 was an example from the ETTm2 dataset with a prediction length of 720. The figure demonstrated that the orange interval was defined by adding and subtracting the values of standard deviation $\sigma$ predicted by the deep probabilistic network. This interval represented the confidence interval corresponding to one standard error above and below the prediction expectation. Furthermore, as was illustrated in this figure, the uncertainty levels increased as the prediction horizon extended. This trend was consistent with expectations, as the model faced greater uncertainty when forecasting events further into the future. Without loss of generality, AdaPRL can improve the interpretability of prediction results by providing an uncertainty estimation for each prediction value for different types of regression tasks.
@@
B.9 Analysis of Out-of-fold Uncertainty Estimation and Prediction Error
@@
To analyze the relationship between out-of-fold uncertainty estimation and algorithm prediction errors, we conducted the following experiments on Movielens-1M and KuaiRec datasets. First, we trained the model using AdaPRL on the training set according to the previous experimental setup. Then, using the auxiliary deep probabilistic network of our model, we made predictions on the test set, obtaining uncertainty estimates $\sigma$ for each sample in the test set, and we obtained point estimation from the main network. Next, we calculated the point estimate test error for the means. Then, we compared the correlation between the standard deviations and the actual test errors. The results were shown in Figures 10 and 11. The evaluation result of the test set was grouped according to the uncertainty estimation predicted by the deep probabilistic network in AdaPRL, and the absolute error distribution of each group was on the y-axis. The diamond signal indicated the average test error of each group. Note that the outliers of KuaiRec dataset had been eliminated for better illustration. As was shown in the figures, when the model’s predicted standard deviation was larger, indicating higher uncertainty in the model’s prediction for these samples, we could indeed observe that the actual test error for that sample tended to be larger, showing an obvious positive correlation between the two. The phenomenon was very similar in both datasets, which clearly demonstrated the effectiveness of the deep probabilistic network of AdaPRL in estimating uncertainties.
Appendix C Experiment Conclusion and Further Study
@@
As we have analyzed in the previous sections, our proposed AdaPRL method has demonstrated notable potential in several key areas. Firstly, it has the capacity to enhance prediction accuracy and ranking ability in various domains, which are crucial for improving the performance of modern deep learning regression models. Secondly, AdaPRL has contributed to increasing generalization capability, enabling the model to perform better on unforeseen data corruption and limited data scenarios. Additionally, it has exhibited improved robustness to noisy data, which is particularly important in real-world applications. Moreover, our method enhances interpretability, allowing practitioners to gain deeper insights into the model’s decision-making process and fostering trust in its predictions. Lastly, the results of the experiments show that AdaPRL can seamlessly incorporate into different regression methods such as Ranksim and Ordinal Entropy, yielding even better accuracy. In conclusion, these combined advantages position our approach as a promising improved solution in the field of regression task learning.
@@
Therefore, we believe this method extends to a broader range of deep learning problems, particularly classification tasks. Uncertainty-based pairwise learning demonstrates exceptional performance improvements and generalization enhancements for regression tasks, leading us to believe that similar benefits apply to classification problems. This represents a research direction with tremendous potential; however, since this paper primarily focuses on regression problems, we do not delve further into this topic here. Instead, we intend to explore this area in our future studies.

### v3

@@ B.3.2 Results
Table 15: Comparison of performance of RankSim, Ordinal Entropy, and their AdaPRL-integrated methods on the KuaiRec, AgeDB, and Jigsaw datasets. Rel.Imp denotes the relative improvement over the corresponding metrics. The experiments were repeated five times and the average results were reported.
B.4 Robustness to Label Noise with AdaPRL
B.4.1 Experiment Setup
Label noise is a common problem in real-world datasets, where mislabeling can occur due to data ambiguity, human error, automated labeling processes, or intrinsic randomness of data, which can significantly degrade the performance of deep neural networks because of its misleading effects. As we mentioned before, AdaPRL could theoretically alleviate aleatoric noise of the data because of the uncertainty estimation module. To prove the theory, we evaluated our method against the standard baseline MSE loss using the KuaiRec dataset under varying levels of Gaussian noise for labels. The noise was introduced into the labels at different levels, where the label noise level $k$ signified that Gaussian noise with zero mean and a standard deviation of $0.2k\times\mathrm{std}(y)$ was added to the labels, and $\mathrm{std}(y)$ denoted the standard deviation of the original labels. This setup allowed us to assess the robustness of our method in the presence of different degrees of label noise. The experiment was repeated five times and the average mean squared error (MSE) of the test data was reported.
B.4.2 Results
The result is presented in Figure 4. As we can see, AdaPRL consistently outperformed the baseline, exhibiting significantly less performance degradation as the proportion of noisy labels increases. Furthermore, as the level of label noise increases, our proposed method exhibits greater relative improvements in the MSE metric compared to the traditional MSE loss. This indicates that our method effectively mitigates the impact of label noise, resulting in more reliable model performance even when training data labels are imperfect.
@@ B.4.2 Results
B.5 Robustness to Unforeseen Data Corruption using AdaPRL
B.5.1 Experiment Setup
Ensuring that deep neural networks maintain performance in the presence of unforeseen data corruptions in the inferring stage is essential for their deployment in real-world applications. Models often encounter various unexpected corruptions or inconsistencies, such as missing data, noise from data collection, and inconsistencies of online and offline features, which could significantly degrade online performance. To simulate the problems, we chose the Movielens-1M and KuaiRec datasets as representative, and the evaluation protocol is as follows. Firstly, we randomly selected 20% of the columns to corrupt. Secondly, for each corruption level $k$, selected $10k\%$ of the validation and test samples. Then, randomly shuffled the values within the selected samples for each column. Finally, trained the model with the original training data and evaluated it on the corrupted test set. Note that in this experiment, our model was trained solely on uncorrupted data but was evaluated on corrupted test sets. Early stoppings were applied on the original validation data, and a coarse-grained grid search over alpha was conducted as mentioned above. The experiment was repeated five times, and the average of mean squared error (MSE) of the test data was reported.
@@ B.5.1 Experiment Setup
B.5.2 Results
As shown in Figure 6, our proposed method AdaPRL achieved evident improvement over the baseline and consistently achieved relative reductions in MSE compared to the L2 loss as the severity of data corruption increased. Furthermore, the relative improvement in MSE gradually increases as the level of data corruption increases, thereby demonstrating the improved resilience of our model to unforeseen test data corruptions.
@@ B.5.2 Results
B.6 Resilience to Reduced Training Data with AdaPRL
B.6.1 Experiment Setup
Although the success of deep learning models is heavily dependent on the scale of training datasets, acquiring large quantities of labeled data is often impractical due to resource constraints. In many real-world situations, models must be effective even with limited training data. In this section, we conducted experiments by creating random subsets of the KuaiRec dataset with varying fractions to simulate scenarios with limited data. Our method was compared against a standard L2 loss baseline under the same conditions. The experiment was repeated five times and the average mean squared error (MSE) of test data was reported.
B.6.2 Results
As shown in Figure 7, our model demonstrates superior performance with significantly less degradation as the number of training samples decreases. This indicates that AdaPRL is more effective in settings where training data is limited, making it a viable solution for applications with constrained data availability. Another conclusion that can be drawn is that AdaPRL outperformed L2 loss in all cases and exhibited the highest relative improvement when the number of training data was at a moderate level.
@@ B.7 Visualization of Time Series Forecasting Prediction Results
To facilitate an instinct comparison among different backbone forecasting models with or without AdaPRL, we present prediction visualizations for two representative datasets in Figure 8. The figure illustrated the prediction results produced by the following models: iTransformer Liu et al. (2024) and Minusformer Liang et al. (2024) with a prediction length of 336. Upon visual inspection of the prediction results across these datasets, both iTransformer and Minusformer trained with AdaPRL demonstrated better accuracy in capturing future series variations, exhibiting superior performance in long-term forecasting compared to using MSE loss, which was consistent with the previous experimental result.
@@ B.8 Visualization of Time Series Forecasting Uncertainty Estimation
By integrating a deep probabilistic network with an arbitrary backbone network, our proposed AdaPRL is capable of generating not only point estimates for each timestamp in long-term time series forecasting but also predicting the associated uncertainties. Figure 9 is an example from the ETTm2 dataset with a prediction length of 720. The figure demonstrates that the orange interval is defined by adding and subtracting the standard deviation $\sigma$ values predicted by the deep probabilistic network. This interval represents the confidence interval corresponding to one standard error above and below the prediction expectation. Furthermore, as is illustrated in this figure, the uncertainty levels increase as the prediction horizon extends. This trend is consistent with expectations, as the model faces greater uncertainty when forecasting events further into the future.
Without loss of generality, AdaPRL enhances the interpretability of prediction results by providing uncertainty estimations for each prediction value across various types of regression tasks. This capability is particularly valuable as it allows practitioners to gauge the reliability of predictions, thereby facilitating more informed decision-making processes. By incorporating uncertainty quantification, AdaPRL not only predicts the target values but also offers insights into the confidence associated with these predictions, which is crucial in applications where the precision and trustworthiness of predictions are important.
@@
B.9 Analysis of Auxiliary Deep Probabilistic Network
B.9.1 Experiment Setup
@@
To better understand the performance of the auxiliary deep probabilistic network as well as the relationship between the out-of-fold uncertainty estimation and prediction errors, we conducted the following experiments on the Movielens-1M and KuaiRec datasets. First, we trained the model using AdaPRL on the training set following former experimental setups. Then, using the auxiliary deep probabilistic network of our model, we made predictions on the test set, obtaining uncertainty estimates $\sigma$ for each sample in the test set, and we obtained the point estimation from the main network. Next, we calculated the point estimate test error for the means and compared the correlation between the standard deviations and the actual test errors.
B.9.2 Results
The results are shown in Figures 10 and 11. The evaluation result of the test set was grouped according to the uncertainty estimation predicted by the deep probabilistic network in AdaPRL, and the absolute error distribution of each group was on the y-axis. The diamond signal indicated the average test error of each group. Note that the outliers of the KuaiRec dataset were eliminated for better illustration. As shown in the figures, when the model’s predicted standard deviation was larger, indicating higher uncertainty in the model’s prediction for these samples, we indeed observed that the actual test error for that sample tended to be larger, showing an obvious positive correlation between the two. The phenomenon is very similar in both datasets, which clearly demonstrates the effectiveness of the deep probabilistic network of AdaPRL in estimating uncertainties.
Dataset Movielens-1M KuaiRec
Training Platform Batch Size time(MSE) time(AdaPRL) Rel.Inc time(MSE) time(AdaPRL) Rel.Inc
CPU Intel Xeon Platinum 8168 256 27.27 49.24 80.6% 496.15 978.54 97.2%
512 16.53 30.96 87.3% 325.36 650.23 99.9%
1024 11.06 20.60 86.3% 258.51 471.60 82.4%
2048 8.63 31.58 265.8% 206.90 449.38 117.2%
4096 8.77 79.25 804.0% 195.70 701.65 258.5%
8192 8.84 150.24 1600.3% 197.84 $1$.05 461.1%
GPU NVIDIA RTX A6000 256 13.64 27.99 105.2% 257.66 527.47 104.7%
512 9.04 15.73 74.0% 174.36 260.12 49.2%
1024 6.12 10.76 75.9% 125.23 170.51 36.2%
2048 6.34 8.21 29.5% 106.69 143.14 34.2%
4096 7.41 8.88 19.9% 113.99 142.30 24.8%
8192 7.76 9.01 16.2% 106.23 123.08 15.9%
Table 16: Training time cost each epoch on Movielens-1M and KuaiRec datasets using MLP as backbone model. The benchmark experiments were repeated 5 times and the minimum of corresponding training time cost were reported. Rel.Inc indicates the relative increase of training time cost of AdaPRL over the baseline MSE loss.
B.10 Theoretical And Experimental Analysis of Extra Computation Costs with AdaPRL
B.10.1 Theoretical Analysis
Earlier experiments have shown that AdaPRL consistently improves the precision of regression tasks as well as the robustness of regression models. Despite all those advantages, it is necessary to conduct an analysis of its computational cost both theoretically and experimentally in order to ensure its efficiency and scalability for large-scale applications.
During the inference stage, if the point estimation is the only concern, as is typically the case in many real-world applications, then the auxiliary network can be ignored and only the original network needs to be deployed; therefore, the inference cost remains exactly the same as the original model, and there will be no extra computational cost during inference at all.
During the training stage, however, extra computation is inevitable for two reasons. Firstly, in any single-output pairwise regression learning (PRL) framework, the difference of the pairwise predictions, the difference of the pairwise labels, as well as their difference are necessary to be calculated, resulting in extra computation cost. Assume the mini-batch size is $B$ then the extra computing cost should be around $O(B^{2})$ for each step. Secondly, since an auxiliary deep probabilistic network is introduced, extra computing is also needed during the training phase in both the network forward and backward processes. If the same backbone is used as the auxiliary network as we did in our experiments, then the training computing task is approximately twice as much as using a single network if the slight difference of the output layer is ignored.
In conclusion, when the batch size in the training phase is small, the additional computational overhead of the auxiliary network is the dominant factor, leading to an approximate doubling of the overall computational cost compared to the original model theoretically. As the batch size increases, however, the extra cost associated with pairwise computations becomes the primary contributor, causing the additional theoretical computational cost to grow proportionally to the square of the batch size.
B.10.2 Experiment Setup
Besides the theoretical analysis, we also conducted multiple comparative benchmark experiments on Movielens-1M and KuaiRec datasets to facilitate a better understanding of the additional time cost of training AdaPRL compared to the baseline loss in actual applications. We trained the models using either CPU or GPU platforms with different batch sizes to assess the overall speed performance differences across hardware platforms. The experiment was run with PyTorch 2.4.1 and CUDA 12.4 environment. For other detailed information of our hardware platform and their corresponding results, please refer to Table 16.
B.10.3 Experimental Results and Analysis
The outcomes of these experiments are depicted in Table 16. As shown in the table, in the context of CPU-based model training, our experimental findings across both datasets aligned with the theoretical analysis. Specifically, for smaller batch sizes ($\leq$1024), the training time for AdaPRL was approximately 1.8 to 2.0 times that of the MSE loss function. Conversely, for larger batch sizes ($\geq$2048), the relative increase in training time for AdaPRL compared to MSE exhibited near-quadratic growth, as the additional computational cost tended to increase roughly with the square of the batch size. These benchmark results confirm the anticipated performance attributes of AdaPRL.
In contrast, GPU-accelerated training yielded different results due to hardware optimizations. For smaller batch sizes (e.g., 256), AdaPRL’s training time was also approximately double that of MSE. However, as batch sizes increased, the relative increment in training time per epoch decreased dramatically from around 105% to around 16%. Consequently, when training regression models for tabular data at moderate batch sizes with GPU, the computational time increase when using AdaPRL instead of MSE is expected to be around 30% and 70%. Given that no additional computational costs are introduced during inference, the computational overhead introduced by AdaPRL is generally considered acceptable in practical industrial scenarios.
Appendix C Conclusion and Further Study
@@
As we have analyzed in previous sections, our proposed AdaPRL method demonstrates notable potential in several key areas. Firstly, AdaPRL has the ability to enhance prediction accuracy and ranking ability in various domains according to our extensive experiments, which are crucial for improving the performance of modern deep learning regression models. Secondly, AdaPRL has contributed to increasing generalization capability, enabling the model to perform better on unforeseen data corruption and limited data scenarios. In addition, it has exhibited improved robustness to noisy data, which is considered particularly important in real-world applications. Moreover, our proposed AdaPRL enhances interpretability because it allows practitioners to gain deeper insights into the model’s decision-making process and fosters trust in its predictions. Lastly, our experiments show that AdaPRL can seamlessly incorporate into different regression methods, such as RankSim and Ordinal Entropy, yielding even better accuracy. In conclusion, these combined advantages position our approach as a promising improved solution in the field of regression tasks.
@@
Therefore, we believe that this method extends to a broader range of deep learning problems, particularly classification tasks. Uncertainty-based pairwise learning demonstrates exceptional performance improvements and generalization enhancements for regression tasks, leading us to believe that similar benefits apply to classification problems. This represents a research direction with great potential; however, since this paper primarily focuses on regression problems, we do not delve further into this topic here. Instead, we intend to explore this area in our future studies.